# DiffInk: Glyph- and Style-Aware Latent Diffusion Transformer for Text to Online Handwriting Generation

### v1

@@ 5 Conclusion
In this paper, we propose DiffInk, a pioneering conditional latent diffusion Transformer framework for TOHG. Our key contributions are: (1) a task-aware variational autoencoder (InkVAE) that incorporates OCR-based and style regularization to learn a semantically structured latent space with disentangled content and style; and (2) a conditional latent diffusion Transformer (InkDiT) that conditions on textual input and reference styles to generate coherent pen trajectories. Built upon these components, DiffInk produces full-line handwritten text with high content accuracy, consistent style, and natural inter-character transitions. Comprehensive experiments demonstrate that DiffInk significantly outperforms existing methods, validating the effectiveness of our approach. This work establishes the effectiveness of the conditional latent diffusion model for complex online handwriting generation and highlights its potential for OCR (see Appendix D.4), personalized digital handwriting applications, and human-computer interaction systems. Limitations and discussions, including extensions to paragraph-level generation and to other languages, are presented in Appendix D.
@@ Appendix A Datasets and Preprocessing
Figure 8: Analysis of text-line length distribution in the Chinese handwriting dataset. Due to dataset constraints, our method supports up to 52 characters per line, which is higher than the average of 26.68 characters reported in (Yin et al., 2013) and sufficient for practical usage. In theory, the character length in our approach is not limited.
@@
For Chinese TOHG, we use the CASIA-OLHWDB 2.0–2.2 (Liu et al., 2011) datasets, which contain over 50,000 online handwritten text lines collected from 1,019 writers. Due to training resource limitations, we select 500 writers based on the average length of their text lines, prioritizing those with longer and more informative content. Among them, 400 writers are used for training and 100 for testing. As shown in Figure 8, we further analyze the text-line length distribution in the training set, where the average length exceeds 30 characters (notably higher than the reported 26.68 characters in prior benchmarks (Yin et al., 2013)) and the maximum length goes beyond 40 characters—sufficient to cover typical text-line writing scenarios in daily life.
@@
We first normalize the $(x,y)$ coordinates of all trajectory points following prior work (Zhang et al., 2017), to address scale inconsistencies caused by variations in device resolution and handwriting styles. This normalization facilitates more stable learning of structural patterns in handwritten text. Subsequently, we apply the Ramer–Douglas–Peucker(RDP) (Douglas & Peucker, 1973) algorithm to each stroke sequence to remove redundant points, thereby reducing the overall trajectory length and improving computational efficiency. The simplification coefficients are empirically selected based on extensive trial experiments: we set them to 0.4 for CASIA-OLHWDB. The average sequence length of Chinese handwriting is reduced by about 500 points after simplification. These values balance preserving handwriting structure with noise reduction, while also shortening the input sequences. For each batch of text-line data, we pad all sequences to the maximum line length. The padding regions are filled with $x=y=0$, while the pen-state penpen is set to the End of Character symbol.
@@ B.1 Details of InkVAE
Recent studies (Tang et al., 2025; Yao et al., 2025; Guo et al., 2025) suggest that a well-structured latent space plays a crucial role in the effectiveness of downstream diffusion-based modeling. To this end, beyond trajectory reconstruction, InkVAE introduces two auxiliary objectives to impose task-specific regularization on the latent space: (1) A style classifier composed of an LSTM (Staudemeyer & Morris, 2019) with attention pooling and a 400-class writer identification head, and (2) a Transformer-based OCR module (hidden size 384) equipped with a 2648-class character classifier and trained with CTC loss to encourage content-aware representation learning. The style classification loss is defined as Equation 6, where $y^{\text{style}}$ is the writer identity label and $p^{\text{style}}$ is the predicted probability. The OCR module is trained using the standard CTC loss (Graves et al., 2006), as shown in Equation 6, to align predicted character sequences with ground-truth transcriptions.
@@ B.4 Evaluation Metrics
The text-line DTW (Berndt & Clifford, 1994) distance is computed over the $(x,y)$ coordinate sequences of each full text line and normalized by the length of the corresponding ground-truth trajectory, following the formulation in Equation 9. Here, $N$ is the number of test samples, $|x_{i}^{\text{gt}}|$ denotes the length of the $i$-th ground-truth trajectory, and $\text{DTW}(\cdot)$ is efficiently computed using the torch-fastdtw library.
@@
The style accuracy is obtained by feeding the generated text lines into a style classifier, predicting their corresponding style categories, and comparing them with the ground-truth labels. To this end, we train a 100-class writer identification model on the training set, which achieves a classification accuracy of 99.9% on the test set, demonstrating its strong discriminative ability for style evaluation. The style classifier consists of the encoder of InkVAE followed by a writer classification head. The generation speed is measured by recording the average time required to generate 2000 characters and taking its reciprocal, which reflects the number of characters the model can generate per second and serves as an indicator of generation efficiency.
@@ Appendix C More Experimental Results
C.1 More Comparisons with existing Methods
We provide additional comparisons with existing approaches, including WLU (Tang & Lian, 2021), SDT (Dai et al., 2023), and OLHWG (Ren et al., 2025). As illustrated in Figure 9 and Figure 10, our method delivers superior text-line generation quality over a broader spectrum of handwriting styles. Notably, character-level generators such as WLU and SDT tend to produce unnatural artifacts when two consecutive characters exhibit large structural differences, leading to visible discontinuities or awkward spacing. In contrast, by modelling entire text lines holistically, DiffInk preserves smoother transitions between adjacent characters and maintains overall visual coherence. It is worth noting that the performance of WLU and SDT heavily relies on pre-trained content or style encoders, which further increases the complexity of deploying and using these methods in practice.
Figure 9: More visual comparisons with SOTA methods. DiffInk generates more natural text-line results, while the red boxes highlight unnatural character connections caused by layout-decoupled methods such as WLU, SDT, and OLHWG. Although these methods produce reasonably good single-character content, their reliance on the current state-of-the-art layout prediction module still leads to noticeable artifacts in character connections.
Figure 10: More visual comparisons with SOTA methods. DiffInk generates more natural text-line results, while the red boxes highlight unnatural character connections caused by layout-decoupled methods such as WLU, SDT, and OLHWG. Although these methods produce reasonably good single-character content, their reliance on the current state-of-the-art layout prediction module still leads to noticeable artifacts in character connections.
C.2 Style Consistency Across Generated Samples.
@@ C.1 More Comparisons with existing Methods
C.3 More Generation Results of DiffInk
In Figure 11, we present additional generation results. We randomly sample eight writers from the test set and generate multiple text lines for each of them. Visual inspection reveals that the generated text lines from the same writer exhibit consistent local stroke patterns and global layout styles, while clear stylistic differences emerge across different writers. These results demonstrate that DiffInk effectively captures inter-writer style variation, and that our lightweight style classification strategy improves the model’s ability to generalize across styles. The generated text lines also maintain high content accuracy and readability, further validating the effectiveness of the proposed DiffInk framework.
Figure 11: More generation results are presented. DiffInk is capable of producing diverse handwritten long text lines, exhibiting high consistency within the same style and clear distinctions across different styles.
@@ D.2 Implementation of DiffInk
While this setup allows for controlled comparison, it remains unrealistic in practice to assume that a reference trajectory is available for every text line. To simulate a more practical and challenging scenario, we adopt a one-shot setting: for a given writer, only a single reference line—or even a short trajectory of fewer than ten characters—is provided as the prompt $X_{\text{ref}}$, with its corresponding text $T_{\text{ref}}$. The user then specifies a target text line $T_{\text{gen}}$, and the model generates the corresponding trajectory $X_{\text{gen}}$. As illustrated in Figure 14, we present qualitative results under this one-shot setting. We observe that, given different target texts $T_{\text{gen}}$, the model is still able to generate text lines with reasonable layouts. Moreover, to further enhance the diversity of generated results, the reference line $X_{\text{ref}}$ can be augmented to simulate variations in layout structures. Furthermore, our approach offers a promising direction for paragraph-level generation. As illustrated in Figure 15, we present an example of a model-generated handwritten paragraph of classical Chinese poetry. By simply concatenating the line-level results, a complete paragraph can be obtained.
@@ D.3 Applying to Other Languages
To validate the generalizability of the proposed method, we directly applied it to English text line generation tasks and conducted experiments on the IAM-OnDB dataset. The data processing pipeline remained consistent with the Chinese setup, including the use of the RDP algorithm (with the redundancy removal parameter set to 0.5) for redundant point removal and normalization of trajectory coordinates. The IAM-OnDB (Liwicki & Bunke, 2005) dataset contains approximately 10,000 handwritten English text line samples produced by 221 writers, featuring substantial cursive writing. We randomly selected data from 25 writers as the test set, with the remaining data used for training. During training, all hyperparameters remained consistent with the Chinese experimental configuration, except for the learning rate, which was reduced to $5\times 10^{-5}$.
@@
D.4 Improving OCR Performance with Synthetic Data
In practical applications, the quantity of handwriting data plays a critical role in improving OCR performance. To assess the impact of our synthesized data, we use the generated text lines to augment the original training set. For recognition, we adopt the same OCR model employed in this work for evaluation—a CNN+Transformer architecture—as the text-line recognition backbone.
Our experimental results on the ICDAR13 dataset (Yin et al., 2013) are presented in Table 4. It can be observed that the proposed generation method significantly improves the performance of the recognition model. Specifically, after augmenting the training set with our synthesized data, the AR and CR metrics increased by 8.5 and 8.6 percentage points, respectively. This demonstrates that our synthesized data can effectively mitigate recognition degradation caused by the long-tail distribution of characters and the limited diversity of handwriting styles, thereby substantially enhancing the representational capacity of existing datasets and further confirming the high content and style quality of our synthetic data.
@@ D.4 Improving OCR Performance with Synthetic Data
CNN+Transformer Real data only (Baseline) 77.2 77.7
+ DiffInk Synth (Ours) 85.7 86.4
D.5 Limitations

### v2

@@ 5 Conclusion
In this paper, we propose DiffInk, a pioneering conditional latent diffusion Transformer framework for TOHG. Our key contributions are: (1) a task-aware variational autoencoder (InkVAE) that incorporates OCR-based and style regularization to learn a semantically structured latent space with disentangled content and style; and (2) a conditional latent diffusion Transformer (InkDiT) that conditions on textual input and reference styles to generate coherent pen trajectories. Built upon these components, DiffInk produces full-line handwritten text with high content accuracy, consistent style, and natural inter-character transitions. Comprehensive experiments demonstrate that DiffInk significantly outperforms existing methods, validating the effectiveness of our approach. This work establishes the effectiveness of the conditional latent diffusion model for complex online handwriting generation and highlights its potential for OCR, personalized digital handwriting applications, and human-computer interaction systems. Limitations and discussions, including extensions to paragraph-level generation and to other languages are presented in Appendix D.
@@ Appendix A Datasets and Preprocessing
Figure 9: Analysis of text-line length distribution in the Chinese handwriting dataset. Due to dataset constraints, our method supports up to 52 characters per line, which is higher than the average of 26.68 characters reported in (Yin et al., 2013) and sufficient for practical usage.
@@
For Chinese TOHG, we use the CASIA-OLHWDB 2.0–2.2 (Liu et al., 2011) datasets, which contain over 50,000 online handwritten text lines collected from 1,019 writers. Due to training resource limitations, we select 500 writers based on the average length of their text lines, prioritizing those with longer and more informative content. Among them, 400 writers are used for training and 100 for testing. As shown in Figure 9, we further analyze the text-line length distribution in the training set, where the average length exceeds 30 characters (higher than the reported 26.68 characters in prior benchmarks (Yin et al., 2013)) and the maximum length goes beyond 40 characters—sufficient to cover typical text-line writing scenarios in daily life.
@@
We first normalize the $(x,y)$ coordinates of all trajectory points following prior work (Zhang et al., 2017), to address scale inconsistencies caused by variations in device resolution and handwriting styles. This normalization facilitates more stable learning of structural patterns in handwritten text. Subsequently, we apply the Ramer–Douglas–Peucker(RDP) (Douglas and Peucker, 1973) algorithm to each stroke sequence to remove redundant points, thereby reducing the overall trajectory length and improving computational efficiency. The simplification coefficients are empirically selected based on extensive trial experiments: we set them to 0.4 for CASIA-OLHWDB. The average sequence length of Chinese handwriting is reduced by about 500 points after simplification. These values balance preserving handwriting structure with noise reduction, while also shortening the input sequences. For each batch of text-line data, we pad all sequences to the maximum line length. The padding regions are filled with $x=y=0$, while the pen-state penpen is set to the End of Character symbol.
@@ B.1 Details of InkVAE
Recent studies (Tang et al., 2025; Yao et al., 2025; Guo et al., 2025) suggest that a well-structured latent space plays a crucial role in the effectiveness of downstream diffusion-based modeling. To this end, beyond trajectory reconstruction, InkVAE introduces two auxiliary objectives to impose task-specific regularization on the latent space: (1) A style classifier composed of an LSTM (Staudemeyer and Morris, 2019) with attention pooling and a 400-class writer identification head, and (2) a Transformer-based OCR module (hidden size 384) equipped with a 2648-class character classifier and trained with CTC loss to encourage content-aware representation learning. The style classification loss and the OCR loss are defined as Equation 6, where the former uses a cross-entropy loss between the writer identity label $y^{\text{style}}$ and the predicted probability $p^{\text{style}}$, and the latter adopts the standard CTC loss (Graves et al., 2006) to align predicted character sequences with ground-truth transcriptions.
@@ B.4 Evaluation Metrics
The text-line DTW (Berndt and Clifford, 1994) distance is computed over the $(x,y)$ coordinate sequences of each full text line and normalized by the length of the corresponding ground-truth trajectory, following the formulation in Equation 9. Here, $N$ is the number of test samples, $|x_{i}^{\text{gt}}|$ denotes the length of the $i$-th ground-truth trajectory, and $\text{DTW}(\cdot)$ is efficiently computed using the torch-fastdtw library.
@@
The style score is obtained by feeding the generated text lines into a style classifier, predicting their corresponding writer identities. To this end, we train a 100-class writer identification model on the testing set, which achieves a classification accuracy of 99.9% on the test set, demonstrating its strong discriminative ability for style evaluation. The style classifier consists of the encoder of InkVAE followed by a writer classification head. Besides, generation efficiency (Avg. char/s) is evaluated as the average number of characters generated per second. All methods are tested with a batch size of 1 on the same NVIDIA RTX A6000 GPU. The average generation speed is computed by dividing the total number of generated characters by the elapsed time, excluding data loading and other preprocessing operations.
@@ Appendix C More Experimental Results
C.1 Style Control and Consistency
Style Consistency
@@ C.1 Style Control and Consistency
Style Control
DiffInk is capable of handling style-mixing inputs, allowing the model to generate handwriting that interpolates between multiple reference styles. As illustrated in the Figure 11, given two style examples and a target text string, the model synthesizes pen-trajectory sequences that simultaneously reflect structural consistency with the textual content and stylistic characteristics drawn from both references.
@@
Figure 11: Multi-style input. DiffInk can take two style references simultaneously and, conditioned on the given text, generate handwriting that blends characteristics of both styles, enabling controllable style manipulation.
C.2 More Comparisons with existing Methods
We provide additional comparisons with existing approaches, including Drawing (Zhang et al., 2017), Deepiminator (Zhao et al., 2020)WLU (Tang and Lian, 2021), SDT (Dai et al., 2023), and OLHWG (Ren et al., 2025). As illustrated in Figure 12 and Figure 13, our method delivers superior text-line generation quality over a broader spectrum of handwriting styles.
Figure 12: More visual comparisons with SOTA methods. DiffInk generates more natural text-line results, while the red boxes highlight unnatural character connections.
Figure 13: More visual comparisons with SOTA methods. DiffInk generates more natural text-line results, while the red boxes highlight unnatural character connections.
C.3 More Results Generated by DiffInk
In Figure 14, we present additional results generated by DiffInk. We randomly sample eight writers from the test set and generate multiple text lines for each of them. Visual inspection reveals that the generated text lines from the same writer exhibit consistent local stroke patterns and global layout styles, while clear stylistic differences emerge across different writers. These results demonstrate that DiffInk effectively captures inter-writer style variation, and that our lightweight style classification strategy improves the model’s ability to generalize across styles. Moreover, the generated text lines also maintain high content accuracy and readability, further validating the effectiveness of the proposed DiffInk framework.
Figure 14: More generation results by DiffInk (Ours). DiffInk is capable of producing diverse handwritten long text lines, exhibiting high consistency within the same style and clear distinctions across different styles.
C.4 DDIM Sampling Steps Evaluation
To evaluate the impact of DDIM sampling steps on generation performance, we conducted an additional experiment. As shown in the table 4, We observe that the performance of DDIM sampling continues to improve as the number of steps increases from 1 to 5. However, when the number of steps exceeds 10, the gains in content and style scores become marginal, while the DTW value increases and the sampling cost grows significantly. Therefore, we adopt 5-step DDIM sampling in this paper as a practical trade-off between generation quality and computational efficiency.
Table 4: DDIM sampling–step analysis of DiffInk, illustrating the trade-off between generation quality and inference speed.
Sample Steps AR% ↑ CR% ↑ Style ↑ DTW ↓ Time Cost
1 36.23 36.39 35.85 1.028 0.2x
2 79.70 80.00 66.31 1.022 0.4x
3 89.93 90.19 73.51 1.026 0.6x
4 93.16 93.39 76.47 1.040 0.8
5 (Ours) 94.38 94.58 77.38 1.052 1x
6 94.86 95.01 77.69 1.058 1.2x
10 95.35 95.52 78.32 1.067 2.0x
15 95.36 95.56 79.10 1.068 3.0x
20 95.27 95.47 79.16 1.076 4.0x
@@ D.3 Applying to Other Languages
To validate the generalizability of the proposed method, we directly applied it to English text line generation tasks and conducted experiments on the IAM-OnDB dataset. The data processing pipeline remained consistent with the Chinese setup, including the use of the RDP algorithm (with the redundancy removal parameter set to 0.5) for redundant point removal and normalization of trajectory coordinates. The IAM-OnDB (Liwicki and Bunke, 2005) dataset contains approximately 10,000 handwritten English text line samples produced by 221 writers, featuring substantial cursive writing. We randomly selected data from 25 writers as the test set, with the remaining data used for training. During training, all hyperparameters remained consistent with the Chinese experimental configuration, except for the learning rate, which was reduced to $5\times 10^{-5}$.
@@
CNN+Transformer Real data only (Baseline) 83.9 84.2
+ DiffInk Synth (Ours) 89.6 90.0
GLRNet (Peng et al., 2021) Real data only (Baseline) 86.2 86.6
+ DiffInk Synth (Ours) 90.9 91.1
D.4 Improving OCR Performance with Synthetic Data
In practical applications, the quantity of handwriting data plays a critical role in improving OCR performance. To assess the impact of our synthesized data, we use the generated text lines to augment the original training set. For recognition, we adopt the same OCR model employed in this work for evaluation—a CNN+Transformer architecture—as the text-line recognition backbone. In addition, we also follow the recognition model design used in GLRNet (Peng et al., 2021) and conduct experiments under the same settings.
Our experimental results on the ICDAR13 dataset (Yin et al., 2013) are presented in Table 5. It can be observed that the proposed generation method significantly improves the performance of the recognition model. Specifically, after augmenting the training set with our synthesized data, the AR and CR metrics increased by 8.5 and 8.6 percentage points, respectively. When using the recognition model from GLRNet (Peng et al., 2021), our synthesized data further yields an average improvement of around 4 percentage points. This demonstrates that our synthesized data can effectively mitigate recognition degradation caused by the long-tail distribution of characters and the limited diversity of handwriting styles, thereby substantially enhancing the representational capacity of existing datasets and further confirming the high content and style quality of our synthetic data.
D.5 Personalized Handwriting Generation with DiffInk
As a text-line handwriting generation model, DiffInk (Ours) is useful not only for producing high-quality synthetic data for OCR systems, but also for broader generative applications such as personalized handwriting services and human–computer interaction. For example, in a personalized handwriting generation scenario: for a given writer, only a single reference line—or even a short trajectory of fewer than ten characters—is provided as the prompt $X_{\text{ref}}$, with its corresponding text $T_{\text{ref}}$. The user then specifies a target text line $T_{\text{gen}}$, and the model generates the corresponding trajectory $X_{\text{gen}}$. As shown in Figure 17, DiffInk yields text lines with coherent layouts under this one-shot setting. To further enhance diversity, the reference trajectory $X_{\text{ref}}$ can be augmented to simulate different structural layouts. Furthermore, our approach offers a promising direction for paragraph-level generation. As illustrated in Figure 18, we present an example of a model-generated handwritten paragraph of classical Chinese poetry. By simply concatenating the line-level results, a complete paragraph can be obtained, further demonstrating that DiffInk can support a broad range of handwriting-related applications.
D.6 Limitations